A 7.00 L tank at 0.543°C is filled with 8.89 g of dinitrogen difluoride gas, 9.41 g sulfur tetrafluoride gas and 9.93 g of chlorine pentafluoride gas. You can assume that the gases behave as ideal gases under these conditions.


Calculate the total pressure in the tank.


1
Expert's answer
2021-12-07T02:09:01-0500

Boron trifluoride = BF3

molar mass = 67.82 g/mol

moles present = 5.42 g x 1 mol / 67.82 g = 0.0799 moles BF3


Sulfur hexafluoride = SF6

molar mass = 146.1 g/mol

moles present = 8.12 mol x 1 mol / 146.1 g = 0.0556 moles SF6


Total moles of gas present = 0.0799 mol + 0.0556 mol = 0.1355 moles of gas

Mole fraction BF3 = 0.0799 / 0.1355 = 0.590

Mole fraction SF6 = 0.0556 / 0.1355 = 0.410


To find total pressure, we can use the ideal gas law PV = nRT

P = pressure = ?

V = volume = 5.00 L

n = moles = 0.1355 moles

R = gas constant = 0.0821 Latm/Kmol

T = temperature in K = 4.43ºC + 273.15 = 277.58K

Solving for P we have...

P = nRT/V = (0.1355)(0.0821)(277.58) / 5.00

P = 0.618 atm = total pressure


Partial pressure BF3 = mole fraction x total pressure = 0.590 x 0.618 atm = 0.365 atm

Partial pressure SF6 = mole fraction x total pressure = 0.410 x 0.618 atm = 0.253 atm